# The Simplest Dark Matter Model at the Edge of Perturbativity

### v1

@@ Abstract
Figure 1: Parameter space for the singlet scalar dark matter model in the multi-TeV scale region. Left: real scalar, right: complex scalar. Purple corresponds to the recent bounds by LZ at 90% C.L. Aalbers et al. (2024). In dashed red we show the values of the coupling that lead to the observed relic abundance at LO in the Higgs portal coupling and in solid red those at NLO. Thus, the regions below the solid red curves would lead to dark matter overabundance today. In this figure we show the results in a renormalization scheme where the direct detection cross section is the same at LO and at NLO. A perturbative treatment of the model is reliable for $\lambda_{\phi H}<15$, see Eq. (9). From the right panel one can see that the entire perturbative parameter space is excluded for the complex scalar case. From the left panel one sees that a three-fold improvement in the sensitivity of direct detection experiments can test the entire perturbative real scalar dark matter TeV-scale window.
@@ III The Singlet Scalar Model at NLO
Relic abundance calculation – To obtain the relic abundance we consider the usual procedure Gondolo and Gelmini (1991) and follow explicitly Ref. Steigman et al. (2012) (which we used also for the tree level results). We take the degrees of freedom in the plasma and their relevant derivatives from Laine and Meyer (2015). We solve the Boltzmann equation numerically taking into account both s-wave and p-wave annihilations and we use as an initial condition for the $\phi$ abundance the thermal number density. We solve the system from $T=m_{\phi}/15$ down to $T=10^{-5}\,m_{\phi}$ using a stiff differential equation solver in Mathematica with a high precision setting. We contrast the final abundance with the measured value as reported by Planck $\Omega_{\rm DM}h^{2}=0.120\pm 0.001$ Aghanim et al. (2020).
@@
Table 1: Summary of the results at LO and at NLO for the three different renormalization schemes as described in the text. DD refers to the bounds from direct detection at 90% C.L. and should be taken as lower bounds for $m_{\phi}$. Pert. refers to the values of the maximum mass that can yield the observed dark matter abundance for the maximum value of the coupling $\lambda_{\phi H}\simeq 15$ for which perturbativity is guaranteed, see Eq. (9).
@@ IV Results
Plugging in this value into the relic density constraint in the three schemes leads to the upper values of the masses quoted in Table 1 for both the real and complex cases. For the complex case, since the direct detection lower bound on the mass is above or very close to these values, one concludes that the complex case is ruled out in its perturbative regime. For the real case as it is well below one concludes that there is still an allowed perturbative regime range that goes from the 20 TeV mass scale quoted above up to possibly about 40-50 TeV. [^2] The fact that in the OS-FO and $\overline{\hbox{MS}}$ schemes the direct detection constraint turn back at already $\sim 34$ TeV and $\sim 26$ TeV respectively, see Fig. 3 (as a result of the fact that the NLO correction is large and has opposite sign with respect to the LO one) does not necessarily mean that we are already lying in the non-perturbative regime, see footnote 1 above, but clearly it is the sign that non-perturbativity is close. Given the dashed purple lines in the figures we can also conclude that if the upper bound on the direct detection cross section improves by a factor of three the full perturbative regime would be ruled out also for the real scalar singlet dark matter case.
@@
Vacuum stability and the sign of the Higgs portal coupling: The range of masses which appears to be still allowed in a perturbative regime for the real case holds only for a positive value of the Higgs portal coupling, $\lambda_{\phi H}>0$. A negative value is excluded in the perturbative regime because the minimum value of the portal coupling that direct detection constraints implies, $|\lambda_{\phi H}|\gtrsim 7$, is too large not to destabilize the vacuum for perturbative values of the $\lambda_{\phi}$ quartic coupling. Vacuum stability requires $\lambda_{\phi H}>-\sqrt{2\lambda_{H}\lambda_{\phi}/3}$ (see e.g. Falkowski et al. (2015)), which, given that the SM quartic coupling $\lambda_{H}$ is equal to $0.13$ would require $\lambda_{\phi}\gg 100$.

### v2

@@ Abstract
Figure 1: Parameter space for the singlet scalar dark matter model in the multi-TeV scale region. Left: real scalar, right: complex scalar. Purple corresponds to the recent bounds by LZ at 90% C.L. Aalbers et al. (2024). In dashed red we show the values of the coupling that lead to the observed relic abundance at LO in the Higgs portal coupling and in solid red those at NLO. Thus, the regions below the solid red curves would lead to dark matter overabundance today. In this figure we show the results in a renormalization scheme where the direct detection cross section is the same at LO and at NLO. A perturbative treatment of the model is expected to be reliable for $\lambda_{\phi H}<15$, see Eq. (9). From the right panel one can see that the entire perturbative parameter space is excluded for the complex scalar case. From the left panel one sees that a three-fold improvement in the sensitivity of direct detection experiments can test the entire perturbative real scalar dark matter TeV-scale window.
@@ III The Singlet Scalar Model at NLO
Relic abundance calculation – To obtain the relic abundance we consider the usual procedure Gondolo and Gelmini (1991) and follow explicitly Ref. Steigman et al. (2012) (which we used also for the tree level results). We take the degrees of freedom in the plasma and their relevant derivatives from Laine and Meyer (2015). We solve the Boltzmann equation numerically taking into account both s-wave and p-wave annihilations and we use as an initial condition for the $\phi$ abundance the thermal number density. We do not take into account thermal effects Biondini et al. (2025) or Sommerfeld enhancement Garcia-Cely et al. (2016) which are both numerically negligible. We solve the system from $T=m_{\phi}/15$ down to $T=10^{-5}\,m_{\phi}$ using a stiff differential equation solver in Mathematica with a high precision setting. We contrast the final abundance with the measured value as reported by Planck $\Omega_{\rm DM}h^{2}=0.120\pm 0.001$ Aghanim et al. (2020).
@@
Table 1: Summary of the results at LO and at NLO for the three different renormalization schemes as described in the text. DD refers to the bounds from direct detection at 90% C.L. and should be taken as lower bounds for $m_{\phi}$. Pert. refers to the values of the maximum mass that can yield the observed dark matter abundance for the value of the coupling $\lambda_{\phi H}\simeq 15$ up to which we expect perturbativity to still hold, see Eq. (9).
@@ IV Results
Plugging in this value into the relic density constraint in the three schemes leads to the upper values of the masses quoted in Table 1 for both the real and complex cases. For the complex case, since the direct detection lower bound on the mass is above or very close to these values, one concludes that the complex case is ruled out in its perturbative regime. For the real case as it is well below one concludes that there is still an allowed perturbative regime range that goes from the 20 TeV mass scale quoted above up to possibly about 40-50 TeV. [^2] The fact that in the OS-FO and $\overline{\hbox{MS}}$ schemes the direct detection constraint turns back at already $\sim 34$ TeV and $\sim 26$ TeV respectively, see Fig. 3 (as a result of the fact that the NLO correction is large and has opposite sign with respect to the LO one) does not necessarily mean that we are already lying in the non-perturbative regime, see footnote 1 above, but clearly it is the sign that non-perturbativity is close. Given the dashed purple lines in the figures we can also conclude that if the upper bound on the direct detection cross section improves by a factor of three the full perturbative regime would be ruled out also for the real scalar singlet dark matter case.
@@
Vacuum stability and the sign of the Higgs portal coupling: The range of masses which appears to be still allowed in a perturbative regime for the real case holds only for a positive value of the Higgs portal coupling, $\lambda_{\phi H}>0$. A negative coupling is excluded in the perturbative regime because the minimum value of the portal coupling that direct detection constraints implies, $|\lambda_{\phi H}|\gtrsim 7$, is too large not to destabilize the vacuum for perturbative values of the $\lambda_{\phi}$ quartic coupling. Vacuum stability requires $\lambda_{\phi H}>-\sqrt{2\lambda_{H}\lambda_{\phi}/3}$ (see e.g. Falkowski et al. (2015)), which, given that the SM quartic coupling $\lambda_{H}$ is equal to $0.13$ would require $\lambda_{\phi}\gg 100$.